OpenStudy (anonymous):

Verify that: y=xe^x is a solution of y''-2y'+y=0. Solve the initial value problem: y(0)=7, y'(0)=4

OpenStudy (lalaly):

\[y'=e^x+xe^x\]\[y''=e^x+e^x+xe^x = 2e^x+xe^x\] substitute in the equation\[2e^x+xe^x-2(e^x+xe^x)+xe^x\]if it equals zero thhen its a solution

OpenStudy (ash2326):

y=xe^x differentiating w.r.t x y'=xe^x+e^x again differentiating w.r.t x y''=xe^x+e^x+e^x y''=xe^x+2e^x now y'=xe^x+e^x i.e y'=y+e^x 2e^x=2y'-2y putting this in y'' equation y''=xe^x+2y'-2y y''=y+2y'-2y y''-2y'-y=0

OpenStudy (ash2326):

y''=2e^x+xe^x on integration we get y'=2e^x+xe^x-e^x+c1 ( c1=constant) y'=e^x+xe^x+c1 now integrating to get y y=e^x+xe^x+c1x-e^x+c2 y=xe^x+c1x+c2 we have to find c1 and c2 y(1)=7 puttiong x=0 in y 7=c2 now y'(0)=4 putting x=0 in y' 4=1+c1 c1=3 so c1=3 and c2=7 y=xe^x+3x+7 y'=e^x+xe^x+3